3 3 A A flower pot from rest to the ground below. If the mass of the pot is m and its initial height is which statement below is correct? A. The KE of the flower pot when it hits the ground does not depend on m. B. The KE of the flower pot when it hits the ground is proportional to h. c. The speed of the flower pot when it hits the ground depends on m. D. The speed of the flower pot when it hits the ground is proportional to h. E. None of the above alternatives are correct
Answer #1


initial energy at top Ei = PE m*g*h

final energy when it hits the pot Ef = KE = (1/2)*m*v^2


from energy conservation

Ef = Ei


PE = KE

m*g*h = (1/2)*m*v^2

v = sqrt(2*g*h)


OPTION B is TRUE


================